An Analysis of the Advancement of the Forensic Science when dealing with Glass

Fragments

Matthew Bartenschlag October 1, 2008

1 1973

1.1 Lawson

Numerical Correlation and Evaluation in the Comparison of Evidentiary Materials

Lawson mentions the use of emission spectroscopy as an identifying characteristic of glass samples. He makes use of ten automobile head lamp glass fragments as the basis of his theory.

1.1.1 Decision Procedure

Lawson developed a technique for ”Sorting the results or getting the smallest numerical index so that the closest match of the parameters involved can be easily evaluated” and ”establish[ing] a threshold or level of reproducibility of the particular technique and material involved.”

1.1.2 Value

This paper appeared to be promising, but the decision procedures created in this document were not further considered.

2 1978

2.1 Seheult

On a Problem in Forensic Science

This paper builds upon Lindley and Evett’s previous papers for using a Bayes approach in determining the likelihood that two glass samples are from the same source. In particular, Seheult defends Lindley’s algorithm using a Neyman-Pearson test.

2.1.1 Value

This paper is not exceptionally valuable to our research because no new decision procedures are provided.

3 1980

3.1 Grove

Interpretation of Forensic Evidence Using a Likelihood Ratio Waiting for Pavan’s paper.

4 1984

4.1 Evett

A Quantitative Theory for Interpreting Transfer Evidence in Criminal Cases

A general approach, based on a Bayesian inference, is presented for providing quantitative interpretation of scientific forensic evidence in cases where material is transferred during the commission of crime. An expression is derived which has a potentially wide applicability in forensic casework and the principles of its use are illustrated by examples.

4.1.1 Value

Evett’s paper is outside the scope of our project, but the use of likelihood ratios continues to be a recurring theme in Evett’s work.

5 1990

5.1 Evett

The Interpretation of Glass Evidence. A Practical Approach

Evett’s 1990 paper attempts to make use of a Bayesian approach even though there are still many considerations are crude.

5.1.1 Value

Evett’s 1990 paper is used as the basis for his 1994 paper. There are graphs that appear in both of these papers, but they are discussed in more detail in Matt Bartenschlag’s research summary of the paper.s